OpenStudy (anonymous):

y^2= 4/5 y + 5/3 Are there two non-real solutions, Two real solutions, or one real solution?

OpenStudy (anonymous):

\[y^2= \frac{4}{5} y + \frac{5}{3}\]

OpenStudy (anonymous):

2 Real